# Domain decomposition-based coupling of Operator Inference Reduced Order Models via the Schwarz alternating method

### v2

@@ 4.1 Static Boundary Conditions
For a first test, we consider the case where $u(\partial\Omega_{T},t)=u(\partial\Omega_{B},t)=0$, $u(\partial\Omega_{L},t)=2$ and $u(\partial\Omega_{R},t)=5\,\,\forall\,t\in(0,1]$. The discontinuity at some of the corners is resolved in favor of the top and bottom boundaries. We establish a baseline in Table 1 by coupling solely FE models in each of the 3 configurations of Figure 2. The wall time, in seconds, was averaged over 5 separate runs. Each subdomain has an overlap of 10 with its non-diagonal neighbors.
@@
The full monolithic simulation, averaged over 5 runs, took 0.94 seconds to run the online portion, and by definition has $\mathbf{E}_{\ell^{2}}^{\text{avg}}=0$.
@@
In Figure 3 and Table 2, we examine the behavior of the OpInf-Schwarz method with respect to $r$, the subdomain overlap, and data by coupling two OpInf models in the vertical orientation of Figure 2(a). Wall clock time results, in seconds, are averaged over 2 individual runs. We will additionally be using a centering strategy on the snapshots, but not scaling the snapshots in any way.
@@
Figure 3: OpInf-Schwarz vs. monolithic simulation. OpInf parameters: $r=6$, Overlap = 10, 30 steps training data.
@@
(a) Fixed parameters: Overlap = 10, $r=6$.
@@
(b) Fixed parameters: 30 steps of training data, $r=6$.
@@
(c) Fixed parameters: 30 steps of training data, Overlap = 10.
@@
From a ROM perspective, Table 2(c) has several interesting results which deserve explanation. First, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ only decreases up to $r=6$, then increases before stalling out for higher values of $r$. This is despite the fact that the projection error of the snapshots onto the basis drops consistently with the basis size. In the row $r=6$, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ and $\mathbf{E}_{\text{proj}}^{\text{avg}}$ have the same order of accuracy, while $\mathbf{E}_{\text{proj}}^{\text{avg}}$ is lower for all further values of $r$. This suggests that the accuracy of the OpInf-Schwarz method is limited by the basis quality up until about $r=6$, and afterwards is most limited by optimization error in the inference step. A basis size of $r=3$ and $r=4$ is necessary to capture 99.9% of the total energy of the snapshots $\Omega_{1}$ and $\Omega_{2}$ respectively.
The second item of note in Table 2(c) is that the average run time goes down by about a third as we increase $r$ from 2 to 6, even though we are increasing the size of the linear systems we are solving. This is due to the coupling of poor quality solutions when $r=2$, which requires additional Schwarz iterations to converge as shown by the mean Schwarz iterations. A similar phenomenon is observed in [3] when coupling subdomain-local intrusive ROMs via the Schwarz alternating method.
The remaining subtables of Table 2 reveal that OpInf-Schwarz is working in the expected manner. In Table 2(b), increasing the size of the overlap decreases the number of Schwarz iterations, and in Table 2(a), increasing the training data decreases the approximation error. Since we are using the same data for both the POD basis and the OpInf training, it is fair to ask whether increasing the data allowance also allows POD to extract useful data for larger values of $r$ – that is, further decreases to $\mathbf{E}_{\text{proj}}^{\text{avg}}$ In our testing, increasing training data had a minor effect on $\mathbf{E}_{\text{proj}}^{\text{avg}}$ compared to increasing $r$, though this would likely become more relevant in more challenging problems. The minimal error we were able to recover was $\mathbf{E}_{\ell^{2}}^{\text{avg}}=2.5\times 10^{-05}$ with the choices of $r=10$, Overlap = 10, and 100 steps of training data.
While the times recorded are certainly not deterministic, a comparison between the FE-FE coupling of Table 1(a) and the minimum and maximum values of Table 2 suggest a speed-up of 8.6 to 15.2 times for fully OpInf-OpInf coupled models as compared to a FOM-FOM coupling on the same spatio-temporal grid, and 2.7 to 4.9 times speed-up for fully OpInf-OpInf coupled models as compared to the monolithic FE simulation. Further speed-ups are possible by introducing the “additive” Schwarz method [41], which computes subdomain solutions in parallel.
@@
We now investigate whether interactions between the subdomain geometry and boundary conditions have an effect on the performance of OpInf-Schwarz. Because the non-zero boundary conditions in this simple case are on the left and right, a vertical overlap splitting the middle has a less complicated gradient than a horizontal overlap. In the more realistic scenario of $r=6$, Overlap = 10, and 30 steps of training data, Table 3 shows that the error and Schwarz iterations are mildly worse for those configurations that impose a more complicated Schwarz boundary.
@@ 4.2 Time-Varying Boundary Conditions
The offset frequencies of the boundary conditions on the left and top boundaries create dynamics which evolve through a significant portion of the time domain, especially in the top left quadrant. These types of problems can be challenging for ROMs to resolve when given insufficient data, though they can usually be fixed by increasing the training data. This is one of the essential choices in ROMs – increasing the training range can often greatly improve accuracy, but data generation in realistic problems can be enormously expensive, especially if it must be generated from a monolithic simulation. Domain decomposition frameworks provide options beyond increasing the training range, and we will now consider a comparison between purely OpInf-Schwarz with a FE-OpInf-Schwarz coupled problem.
4 Squares Configuration
@@
The FE assimilation in Figure 4(b) is not perfect if one looks closely along $x=0$ or $y=0$, but it shows a cohesive solution very similar to the monolithic solution in 4(d), while the solution in Figure 4(a) is clearly failing and is in fact about to diverge entirely. The fully OpInf model with a higher data allowance in Figure 4(c) is comparable to Figure 4(b), and both fail to resolve some aspects of the monolithic simulation.
(a) OpInf solution in all 4 subdomains, 40 steps of training data
(b) FE solution in top left $\Omega_{3}$ only, 40 steps of training data for OpInf portions
(c) OpInf solution in all 4 subdomains, 65 steps of training data
(d) Monolithic FOM (Finite Element Simulation)
@@
One interesting feature of Figure 4(a) is that the Schwarz communication pattern is clearly evident. The solution is clearly incorrect in the top left subdomain, $\Omega_{3}$. The error in $\Omega_{3}$ subsequently propagates to neighboring $\Omega_{1}$ and $\Omega_{4}$, while the bottom right subdomain $\Omega_{2}$ remains closest to an accurate solution.
Following from the results in Table 4, a purely OpInf coupled Schwarz model would require 65 total timesteps (an additional 25 steps of training data) to exceed the average relative accuracy of the OpInf-Schwarz with one FE subdomain and 40 steps of training data (as is displayed in Figure 4(b)).
@@
Table 4: OpInf-Schwarz errors for various training data ranges. $r=6,$ Overlap = 10.
Of course, this stability comes at a price: over two separate runs, the average time taken for the 4 coupled OpInf models in Table 4(a) was 0.73 seconds, while the average for the substitution of one FE model in $\Omega_{3}$ in Table 4(b) is 3.15 seconds, a 4.3 times slowdown over a purely OpInf coupled model. The Schwarz framework gives us the opportunity to more evenly weigh the opportunity cost for amount of data required vs. computational speed for specific problems.
In comparison, a completely FE coupled model with the same settings has $\mathbf{E}_{\ell^{2}}^{\text{avg}}=\mathcal{O}(10^{-14})$, but, over three runs, took an average of 7.9 seconds. A monolithic FE sim takes approximately 0.8 seconds. The “All OpInf” model is therefore about 9 times faster, and the 3 OpInf 1 FE model about 2.4 times faster than the fully FE coupled model. Comparing against the monolithic FOM, the “All OpInf” model is 1.1 times faster and the 3 OpInf 1 FE is 3.9 times slower.
@@
We summarize our timing results for the time-varying boundary problem in the Pareto plot of Figure 5 using data from Table 4 for the “4 Squares” labelled results, and additional results for the “Vertical” labelled results produced in the vertical configuration from Figure 2(a). In particular, “Vertical FE-OpInf” was produced with the FE model in the left subdomain, and “Horizontal FE-OpInf” was produced with the FE subdomain as the top subdomain, both advantageous DDs for resolving the more complicated dynamics in the top left portion of $\Omega$. Fixed parameters are $r=6$ and Overlap = 10. This chart demonstrates the significant impact that subdomain geometry can have when interacting with more complicated boundary issues or solution features. The vertical OpInf-OpInf is more accurate for smaller data allowances, but it is overtaken in accuracy by the 3 OpInf 1 FE model for higher data allowances. All of the models with a FE component have similar time costs, but the horizontal and 4 square configurations make better use of additional data than the vertical configuration does.
Figure 5: Pareto plot for time-varying BCs, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ vs Time (s). Fixed parameters: $r=6$, Overlap $=10$. Plotted points are “Data” as it ranges over 30, 40, 50, and 60 as in Table 4. Values in bottom left are preferred.
@@ 5 Conclusions and Future Work
A major impetus driving the development of the OpInf-Schwarz method is the desire for a minimally intrusive methodology for constructing ROMs that can interface modularly with existing high performance codes. A truly data-driven and modular domain decomposition-based ROM has the potential to greatly reduce computational expenses incurred by meshing and re-meshing complicated 3D objects, and therefore in performing long run-time and/or multi-query simulations.
The preliminary investigation presented herein has assessed various characteristics of the OpInf-Schwarz framework in the context of a 2D heat equation. Our results have shown that the OpInf-Schwarz method is capable of coupling between both OpInf and FE subdomains to recover accurate solutions, and that a purely OpInf formulation can run faster than a monolithic simulation. We have also demonstrated that OpInf-Schwarz is very flexible, and can be tuned in several ways to the problem at hand.
At the same time, this work has uncovered several challenges that are worth investigating in the future. First, the calculated average relative pointwise error presented in Section 4.1 stagnates at around $\mathcal{O}(10^{-5})$ and does not decrease further despite a increase to the POD basis size and training data. While this level of error is acceptable for many applications, it would be ideal for the method to be truly convergent to the data source. While it seems likely that the issue lies with the optimization error of the regression in problem (9) given that the projection errors reported in Table 2(c) continue to decrease, a comparison with a standard intrusive projection ROM would confirm whether the cause is optimization error or errors induced by the coupling process. It may be possible to reduce optimization errors by employing a technique called re-projection [25], and/or by improving the variational consistency of the model [11].
@@
Currently, the learned boundary operator $\hat{\mathbf{B}}$ scales with the size of the boundary in our formulation but only transmits boundary information weakly at a ROM level into the interior of each subdomain. It is likely that a more sophisticated boundary enforcement strategy is needed. One promising approach is based on centering the reduced POD basis [11] (similar to what is commonly done for conventional intrusive ROMs [12]), as opposed to centering the snapshots, as was done in this paper. Reducing the computational complexity of our method requires reducing the number of columns of $\hat{\mathbf{B}}$, which can be large for multi-dimensional problems with many boundary nodes. To mitigate the cost associated with evaluating $\hat{\mathbf{B}}\mathbf{g}$, it may be possible to perform a separate dimension reduction of this term. Finally, as noted earlier, the overall computational complexity of OpInf-Schwarz can be improved by introducing “additive” Schwarz, characterized by parallel subdomain solves with asynchronous boundary condition communication [41].
The last major point is on extending the Opinf-Schwarz method to more challenging and realistic problems, as we expect the greatest utility of this model to be shown in more complicated situations. The heat equation can be revealing for a new model, but extending the model to 3D, non-linear, convection dominated, and parametric problems [39] is a priority, as are implementing strategies to choose optimal domain decomposition and model assignment to enable on-the-fly ROM-FOM switching for maximal performance.
@@ Acknowledgements
The authors would like to thank Dr. Shane McQuarrie for numerous helpful technical discussions in the area of operator inference, and for providing support on the OpInf library, used to generate the results presented in this paper.

### v3

@@ 4.1 Static Boundary Conditions
For a first test, we consider the case where $u(\partial\Omega_{T},t)=u(\partial\Omega_{B},t)=0$, $u(\partial\Omega_{L},t)=2$ and $u(\partial\Omega_{R},t)=5\,\,\forall\,t\in(0,1]$. The discontinuity at some of the corners is resolved in favor of the top and bottom boundaries. We establish a baseline in Table 1 by coupling solely FE models in each of the 3 configurations of Figure 2. The wall time, in seconds, was averaged over 5 separate runs. Each subdomain has an overlap of 10 with its non-diagonal neighbors. The full monolithic simulation, averaged over 5 runs, takes 0.94 seconds to run the online portion.
@@
In Figure 3 and Table 2, we examine the behavior of the OpInf-Schwarz method with respect to $r$, the subdomain overlap, and data by coupling two OpInf models in the vertical orientation of Figure 2(a). Wall clock time results, in seconds, are averaged over 2 individual runs.
@@
Figure 3: OpInf-Schwarz vs. monolithic simulation. OpInf parameters: $r=6$, Overlap = 10, Data = 30.
@@
(a) Fixed parameters: Overlap = 10 a $r=6$.
@@
(b) Fixed parameters: Data = 30, $r=6$.
@@
(c) Fixed parameters: Data = 30, Overlap = 10.
@@
From a ROM perspective, Table 2(c) has several interesting results which deserve explanation. First, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ only decreases up to $r=6$, then increases before stalling out for higher values of $r$. This is despite the fact that the projection error of the snapshots onto the basis drops consistently with the basis size. A basis size of $r=3$ and $r=4$ is necessary to capture 99.9% of the total energy of the snapshots $\Omega_{1}$ and $\Omega_{2}$ respectively.
In the row $r=6$, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ and $\mathbf{E}_{\text{proj}}^{\text{avg}}$ have the same order of accuracy, while $\mathbf{E}_{\text{proj}}^{\text{avg}}$ is lower for all further values of $r$. This suggests that the accuracy of the OpInf-Schwarz method is limited by the basis quality up until $r=6$, and afterwards is most limited by error in the inferred operators. This could be due to several reasons, such as data quality, our regularization strategy affecting accuracy, or perhaps the Schwarz coupling method interacting with the ROM in an unexpected way.
Another item of note in Table 2(c) is that the average run time decreases by about a factor of three as we increase $r$ from 2 to 6, even though we are increasing the size of the linear systems we are solving. This is due to the coupling of poor quality solutions when $r=2$, which requires additional Schwarz iterations to converge as shown by the mean Schwarz iterations. A similar phenomenon is observed in [3] when coupling subdomain-local intrusive ROMs via the Schwarz alternating method.
The remaining subtables of Table 2 reveal that OpInf-Schwarz performs as anticipated. In Table 2(b), increasing the size of the overlap decreases the number of Schwarz iterations, and in Table 2(a), increasing the training data decreases the approximation error. Since we are using the same data for both the POD basis and the OpInf training, it is fair to ask whether increasing the data allowance also allows POD to extract useful data for larger values of $r$—that is, further decreases to $\mathbf{E}_{\text{proj}}^{\text{avg}}$ In our testing, increasing training data had a minor effect on $\mathbf{E}_{\text{proj}}^{\text{avg}}$ compared to increasing $r$, though this would likely become more relevant in more challenging problems. The minimal error we were able to recover was $\mathbf{E}_{\ell^{2}}^{\text{avg}}=2.5\times 10^{-05}$ with the choices of $r=10$, Overlap = 10, and 100 steps of training data.
While the times recorded are certainly not deterministic, a comparison between the FE-FE coupling of Table 1(a) and the minimum and maximum values of Table 2 suggest a speed-up of 8.6 to 15.2 times for fully OpInf-OpInf coupled models as compared to a FOM-FOM coupling on the same spatio-temporal grid, and 2.7 to 4.9 times speed-up for fully OpInf-OpInf coupled models as compared to the monolithic FE simulation. Further speed-ups may be possible by introducing the “additive” Schwarz method [39], which computes subdomain solutions in parallel.
@@
We now investigate whether interactions between the subdomain geometry and boundary conditions have an effect on the performance of OpInf-Schwarz. Because the non-zero boundary conditions in this simple case are on the left and right, a vertical overlap splitting the middle has a less complicated gradient than a horizontal overlap. In the scenario of $r=6$, Overlap = 10, and 30 steps of training data, Table 3 shows that the error and Schwarz iterations are mildly worse for those configurations that impose a more complicated Schwarz boundary.
@@ 4.2 Time-Varying Boundary Conditions
The offset frequencies of the boundary conditions on the left and top boundaries create dynamics which evolve through a significant portion of the time domain, especially in the top left quadrant. These types of problems can be challenging for ROMs to resolve when given insufficient data, though they can usually be fixed by increasing the training data. This is one of the essential choices in ROMs – increasing the training range can often greatly improve accuracy, but data generation in realistic problems can be enormously expensive, especially if it must be generated from a monolithic simulation. Domain decomposition frameworks provide options beyond increasing the training range, and we now consider a comparison between purely OpInf-Schwarz with a FE-OpInf-Schwarz coupled problem.
Four Squares Configuration
@@
The FE assimilation in Figure 4(b) is not perfect if one looks closely along $x=0$ or $y=0$, but it shows a cohesive solution very similar to the monolithic solution in Figure 4(d), while the solution in Figure 4(a) is clearly failing and is in fact about to diverge entirely. The All OpInf model with a higher data allowance in Figure 4(c) is comparable to Figure 4(b).
(a) OpInf solution in all 4 subdomains, 40 steps of training data in all subdomains.
(b) FE solution in top left $\Omega_{3}$ only, 40 steps of training data for OpInf portions.
(c) OpInf solution in all 4 subdomains, 65 steps of training data in all subdomains.
(d) Monolithic FOM (Finite Element Simulation).
@@
One interesting feature of Figure 4(a) is that the Schwarz communication pattern is clearly evident. The solution is completely incorrect in the top left subdomain, $\Omega_{3}$. The error in $\Omega_{3}$ subsequently propagates to neighboring $\Omega_{1}$ and $\Omega_{4}$, while the bottom right subdomain $\Omega_{2}$ remains closest to an accurate solution.
@@
Table 4: OpInf-Schwarz errors for various training data ranges with $r=6$ and Overlap = 10.
From Table 4, it can be inferred that a purely OpInf coupled Schwarz model would require 65 total timesteps (an additional 25 steps of training data) to exceed the average relative accuracy of the OpInf-Schwarz with one FE subdomain and 40 steps of training data (as is displayed in Figure 4(b)). However, over two separate runs, the average time taken for the 4 coupled OpInf models in Table 4(a) was 0.73 seconds, while the average for the substitution of one FE model in $\Omega_{3}$ in Table 4(b) is 3.15 seconds, a 4.3 times slowdown over a purely OpInf coupled model. The relative ease of model switching in the Schwarz framework lets one choose between data generation costs and monolithic FOM costs.
In comparison, a completely FE coupled model with the same settings has $\mathbf{E}_{\ell^{2}}^{\text{avg}}=\mathcal{O}(10^{-14})$, but, over three runs, takes an average of 7.9 seconds. A monolithic FE sim takes approximately 0.8 seconds. The “All OpInf” model is therefore about 9 times faster, and the 3 OpInf 1 FE model about 2.4 times faster than the fully FE coupled model. Comparing against the monolithic FOM, the “All OpInf” model is 1.1 times faster and the 3 OpInf 1 FE is 3.9 times slower.
@@
Figure 5: Pareto plot for time-varying BCs, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ vs Time (s). Fixed parameters: $r=6$, Overlap $=10$. Plotted points are “Data” as it ranges over 30, 40, 50, and 60 marked by $\times$, $\blacklozenge$, $\boldsymbol{+}$, and $\bullet$ respectively as in Table 4. Values in bottom left are preferred.
We summarize our timing results for the time-varying boundary problem in the Pareto plot of Figure 5 using data from Table 4 for the “Four Squares” labelled results, and additional results for the “Vertical” labelled results produced in the vertical configuration from Figure 2(a). In particular, “Vertical FE-OpInf” was produced with the FE model in the left subdomain, and “Horizontal FE-OpInf” was produced with the FE subdomain as the top subdomain, both advantageous DDs for resolving the more complicated dynamics in the top left portion of $\Omega$. Fixed parameters are $r=6$ and Overlap = 10. This chart demonstrates the significant impact that subdomain geometry can have when interacting with more complicated boundary issues or solution features. The vertical OpInf-OpInf is more accurate for smaller data allowances, but it is overtaken in accuracy by the 3 OpInf 1 FE model for higher data allowances. All of the models with a FE component have similar time costs, but the horizontal and four square configurations make better use of additional data than the vertical configuration does.
@@ 5 Conclusions and Future Work
The major impetus driving the development of the OpInf-Schwarz method is to achieve a minimally intrusive methodology for constructing ROMs that can interface modularly with existing high performance codes. A truly data-driven and modular domain decomposition-based ROM has the potential to greatly reduce computational expenses incurred by meshing and re-meshing complicated 3D objects, and therefore in performing long run-time and/or multi-query simulations.
The preliminary investigation presented herein has assessed various characteristics of the OpInf-Schwarz framework in the context of a 2D heat equation. Our results have shown that the OpInf-Schwarz method is capable of coupling OpInf and FOM subdomains to recover accurate solutions, and that a purely OpInf formulation can run faster than a monolithic simulation. We have also demonstrated that OpInf-Schwarz is very flexible, and can be tuned in several ways to the problem at hand (e.g., by choosing a different DD or ROM/FOM assignment to the subdomains).
At the same time, this work has uncovered several challenges that are worth investigating in the future. First, the calculated average relative pointwise error presented in Section 4.1 stagnates at around $\mathcal{O}(10^{-5})$ and does not decrease further despite a increase to the POD basis size and training data. While this level of error is acceptable for many applications, it would be ideal for the method to be truly convergent to the data source. While it seems likely that the issue lies with the error of the inferred operators from the regression problem (14) given that the projection errors reported in Table 2(c) continue to decrease, a comparison with a standard intrusive projection ROM would confirm whether the cause is error in the inferred operators or errors induced by the coupling process. It may be possible to reduce optimization errors by employing a technique called re-projection [26].
@@
Currently, the size of the learned boundary operator $\hat{\mathbf{B}}$ scales with the size of the boundary in our formulation, but considering that even in this simple 2D case the number of boundary nodes is already 20 or more times larger than the ROM dimension, it is worth questioning whether all of this boundary information is being used in a meaningful way. Reducing the computational complexity of our method requires reducing the number of columns of $\hat{\mathbf{B}}$, which can be large for multi-dimensional problems with many boundary nodes. To mitigate the cost associated with evaluating $\hat{\mathbf{B}}\mathbf{g}$, it may be possible to perform a separate dimension reduction of this term. Finally, as noted earlier, the overall computational complexity of OpInf-Schwarz can be improved by introducing “additive” Schwarz, characterized by parallel subdomain solves with asynchronous boundary condition communication [39].
The last major point is on extending the Opinf-Schwarz method to more challenging and realistic problems, as we expect the greatest utility and time savings of this model to be shown in more complicated situations. The heat equation is a standard prototype for model reduction methods, but extending the model to 3D, non-linear, convection dominated, and parametric problems [22, 38, 40] is a priority, as are implementing strategies to choose optimal domain decomposition and model assignment to enable on-the-fly ROM-FOM switching for maximal performance.
@@ Acknowledgements
The authors would like to thank Dr. Shane McQuarrie for numerous helpful technical discussions in the area of operator inference, and for providing support on the OpInf library.